# MALADY: Multiclass Active Learning with Auction Dynamics on Graphs

### v1

@@ 3.3 Multiclass Active Learning Procedure
We now describe our overall Multiclass Active Learning with Auction Dynamics on graphs (MALADY) procedure. Given an initially labeled set $\mathcal{L}$ with corresponding labels $\{y(x_{i})\}_{x_{i}\in\mathcal{L}}$, we sequentially select query points that maximize our proposed acquisition function (27), $\mathcal{A}(x)=1-\mathcal{M}(x)$. We summarize our method as follows:
@@
    Using the input data, construct a similarity graph using a chosen similarity function such as (1).
@@
    With a current labeled set $\mathcal{L}$, run the underling semi-supervised classifier (Algorithm 1) to get the optimal price and incentive vectors.
@@
    Using the acquisition function (27), select the (next) query point as the maximizer, $x^{\ast}=\operatorname*{arg\,max}_{x\in\mathcal{U}}\mathcal{A}(x)$.
@@
    Query for the label $y^{\ast}$ of query point $x^{\ast}$, and update labeled set $\mathcal{L}\leftarrow\mathcal{L}\cup\{x^{\ast}\}$.
@@
Input : Vertex set $V$, Weight matrix $W$, initial partition $\mathbf{V}$, lower bound $\mathbf{B}$, upper bound $\mathbf{U}$, time step $\delta t$, number of steps $s$, auction error tolerance $\epsilon_{min}$, epsilon scaling factor $\alpha$, initial epsilon $\epsilon_{0}$, graph G = (X, W), initial labeled set $\mathcal{L}_{0}$ , acquisition function $\mathcal{A}(x)$, and budget $T$.
Result: Labeled set $\mathcal{L}$
Initialization: $\mathcal{L}=\mathcal{L}_{0}$;
while $|\mathcal{L}|\leq T$ do
      Run Algorithm 1 (Semi-supervised classification method):  $(\mathbf{V}^{s},\mathbf{p},\mathbf{t})=\text{Semi-supervised classification method}(V,W,\mathbf{V},B,U,\delta t,s,\epsilon_{min})$;
      Pick $x^{\ast}\in\mathcal{U}=V\setminus\mathcal{L}\text{ which has highest acquisition value }\mathcal{A}(x)$;
      $\mathcal{L}\leftarrow\mathcal{L}\cup\{x^{\ast}\}$;
end while
return ${\mathcal{L}}$
@@ 4.1 Data sets
For the computational experiments in this paper, we have used four real-world data sets. The data sets are the following:
  • The Landsat data set [2] is a data set of 6435 elements containing multi-spectral values of pixels in 3 × 3 neighbourhoods in a satellite image. Here, the classification is associated with the central pixel and there are 6 classes. Each of the elements of the data contains the pixel values in the four spectral bands of each pixel in the 3 × 3 neighbourhood. Each feature vector consists of 36 values. The six classes are red soil, cotton crop, grey soil, damp grey soil, soil with vegetation stubble, and very damp grey soil.
  • The USPS data set [3] is a data set of 9,298 square grayscale images which contain 16×16 pixels, using a 4-bit grey scale system with 256 gray levels per pixel. The images were obtained by automatically scanning the envelopes by the United States Postal Service (USPS) in Buffalo, New York. The grayscale images are centered, normalized and show a broad range of font styles. There are 10 classes; each data element is an image of a handwritten digit ”0” through ”9”.
  • The Coil-20 data set [18] is a multiclass data set of 1440 normalized images of 20 objects. The objects were placed on motornized turn table. With a fixed camera, the turn table was rotated through 360 degrees to capture different poses every 5 degrees. There are total of 72 images per object.
@@
Opt-Digits [41] 5620 64 10
@@ 4.2 Comparison to other methods
In this section, we compare our proposed method to other active learning acquisition functions in various graph-based semi-supervised methods. The results of the experiments and the comparison methods are shown in figure 2. In particular, for all data sets, we compare our proposed method, MALADY, to the following methods.
@@
For a fair comparison, we used Algorithm 1 as the underlying semi-supervised method for Random and V-OPT methods. Now, we provide a brief overview of the methods that we include in our numerical comparisons: In the V-OPT algorithm [25], the authors analyze the probability distribution of the unlabeled vertices conditioned on the label information, which follows a multivariate normal distribution with the mean corresponding to the harmonic solution across the field. The nodes are then selected for querying in such a way that the total variance of the distribution on the unlabeled data, as well as the expected prediction is minimized. In [38], the authors design an acquisition function that measures uncertainty in Poisson reweighed Laplace learning algorithm (PWLL). In particular, the authors control the exploration versus exploitation tradeoff in the active learning process by introducing a diagonal perturbation in PWLL which produces exponential localization of solutions.
For the implementation of MALADY, V-OPT, and Random, we utilized the GraphLearning python package [12] which has both graph learning and active learning framework. For the implementation of Unc. (Norm) [38], we utilize the code obtained from the author’s repository.
@@ 4.3 Hyperparameters Selection
In this section, we outline the hyperparameters that we have selected for the experiments. In particular, to make the weight matrix sparse, we compute the weight matrix using a $N_{n}-$ nearest neighbor graph. The number of nearest neighbors is one of the parameters to tune. For our semi-supervised learning framework, we have several parameters. For the computational experiments, the number of steps $s$, auction error tolerance $\epsilon_{min}$, epsilon scaling factor $\alpha$, and the initial epsilon $\epsilon_{0}$ are set to 100, $61e^{-}$ 6, 4, and $71e^{-}$ 7, respectively. We also incorporate class size constraints for the classification which are controlled by upper bound $U$ and lower bound $B$. In computational experiments, we use exact class size constraints. Overall, we outlined all the parameters used for the computational experiments in the Supplementary Information.
@@ 4.4 Performance and Discussion
The data sets that we used for our computational experiments are detailed in Section 4.1 and Table 1. For all data sets, we consider accuracy as the main evaluation metric. For each data set, we used five initial labels per class. The initial labeled set has total of $5\times C$ labeled points where $C$ is total number of labels. We sequentially query 500 additional points using the Algorithm 2. In Figure 2, we show the accuracy performance of each acquisition function averaged over the 10 trails. For all data sets, MALADY has outperformed the other comparison methods.
a) [Uncaptioned image] b) [Uncaptioned image]
Figure 2: Accuracy v.s. # of labeled points for each dataset: (a) Optdigits (b) USPS (c) Satellite (d) COIL-20
c) [Uncaptioned image] d) [Uncaptioned image]
5 Conclusion and Future work
In this paper, we introduce the Multiclass Active Learning with Auction Dynamics on Graphs (MALADY) algorithm which integrates active learning with auction dynamics techniques for semi-supervised learning and data classification, and ensures both exploration and exploitation in the active learning process. The proposed method allows one to obtain accurate results even in the case of very small labeled sets, a common scenario for many applications due to the fact that it is costly in time and expensive to obtain labeled data in many cases. This is in part aided by the active learning of the proposed technique which judiciously selects a limited number of unlabeled data points for labeling at each iteration. The process is aided by incorporating a novel acquisition function derived using the dual variable of an auction algorithm. Moreover, the algorithm allows one to incorporate class size constraints for the data classification task, which improves accuracy even further. Overall, the proposed procedure is a powerful approach for an important task of machine learning.
Declarations:
Conflict of Interest: The authors declare no conflict of interest.
Data availability: The links to all data generated or analyzed during this study is included in this published article.
Acknowledgements
This work is supported in part by NSF grant DMS-2052983.

### v2

@@ 3.3 Multiclass Active Learning Procedure
We now describe our proposed Multiclass Active Learning with Auction DYnamics on graphs (MALADY) procedure. Given an initially labeled set $\mathcal{L}_{0}$ with corresponding labels $\{y(x_{i})\}_{x_{i}\in\mathcal{L}}$, we sequentially select query points that maximize our proposed acquisition function (26), $\mathcal{A}(x)=1-\mathcal{M}(x)$. We summarize our method as follows: Given input data and an initially labeled subset $\mathcal{L}_{0}$, our proposed method consists of the following steps:
@@
    Construct a similarity graph from the data using a chosen similarity function such as (1).
@@
    With a current labeled set $\mathcal{L}$, run the underlying semi-supervised classifier (Algorithm 1) to get the optimal price and incentive vectors. For the first iteration, $\mathcal{L}=\mathcal{L}_{0}$.
@@
    Using the acquisition function (26), select the (next) query point as its maximizer: $x^{\ast}=\operatorname*{arg\,max}_{x\in\mathcal{U}}\mathcal{A}(x)$.
@@
    Query for the label $y^{\ast}$ of the query point $x^{\ast}$ and update the labeled set $\mathcal{L}\leftarrow\mathcal{L}\cup\{x^{\ast}\}$.
@@
  6. 6.
    With a current labeled set $\mathcal{L}$, run the underlying semi-supervised classifier (Algorithm 1) to get the partition of the vertex set, i.e., the data.
  7. 7.
    Evaluate the classifier’s performance by calculating the chosen evaluation metric on the unlabeled set.
The flowchart of MALADY is shown in Figure 3, and the algorithm is also described in detail as Algorithm 2.
@@ 4.1 Data sets
For the computational experiments in this paper, we have used six benchmark data sets, including two hyperspectral data sets. The data sets are the following:
  • The Landsat data set [47] is a data set of 6435 elements containing multi-spectral values of pixels in 3x3 neighborhoods in a satellite image. For this data set, the classification is associated with the central pixel and there are six classes. Each of the data elements contains the pixel values in the four spectral bands of each pixel in the 3×3 neighborhood. Therefore, each feature vector consists of 36 values. The six classes are red soil, cotton crop, grey soil, damp grey soil, soil with vegetation stubble, and very damp grey soil.
  • The USPS data set [48] is a data set of 9,298 square grayscale images that contain 16×16 pixels, using a 4-bit grey scale system with 256 gray levels per pixel. The images were obtained by automatically scanning envelopes by the United States Postal Service in Buffalo, New York. The grayscale images are centered and normalized and show various font styles. There are 10 classes; each data element is an image of a handwritten digit “0” through “9”.
  • The Jasper Ridge data set [49] is one of the most widely used hyperspectral data sets in the hyperspectral image classification task. It consists of an $512\times 614$ image with 198 channels. For our experiments, we only consider a subimage of the image consisting of $100\times 100$ pixels with 198 channels.
  • The Urban data set [50] is one of the most widely used hyperspectral data sets in hyperspectral applications. It consists of an $307\times 307$ image with 162 channels. The version of the data set we use constains four endmembers corresponding to asphalt, grass, tree and roof. The spectral resolution is 10 nm, and each pixel corresponds to an area of two square meters. The data set was recorded in 1995 by the Hyperspectral Digital Imagery Collection Experiment in Copperas Cove, Texas.
  • The COIL-20 data set [51] is a multiclass data set of 1440 normalized images of 20 objects. The objects were placed on a motornized turn table. With a fixed camera, the turn table was rotated through 360 degrees to capture different poses every 5 degrees. There are a total of 72 images per object.
@@
Jasper Ridge [49] 10000 198 4
Urban [50] 94249 162 4
a) Refer to caption
b) Refer to caption
c) Refer to caption
d) Refer to caption
e) Refer to caption
f) Refer to caption
Figure 4: Accuracy v.s. # of labeled points for each dataset: (a) Optdigits (b) USPS (c) Satellite (d) COIL-20 (e) Urban (f) Jasper Ridge
@@ 4.2 Comparison to other methods
In this section, we compare our proposed technique to those using other active learning acquisition functions in various graph-based semi-supervised methods. The results of the experiments and the comparison methods are shown in Figure 2. In particular, for all data sets, we compare our proposed method, MALADY, to the following methods.
@@
  • Uncertainty sampling (Unc-Laplace) [1]
  • k-means++ [53]
For a fair comparison, we used Algorithm 1 as the underlying semi-supervised method for the Random, V-OPT and k-means++ algorithms. For minimum norm uncertainty sampling, we used Poisson reweighted Laplace learning [54] as the underlying method. For Unc-Laplace, the underlying classifier is Laplace learning [55].
Now, we provide a brief overview of the methods that we include in our numerical comparisons. In the V-Opt algorithm [5], the authors analyze the probability distribution of the unlabeled vertices conditioned on the label information, which follows a multivariate normal distribution with the mean corresponding to the harmonic solution across the field. The nodes are then selected for querying such that the total variance of the distribution on the unlabeled data, as well as the expected prediction, is minimized. In minimum norm uncertainty sampling [4], the authors design an acquisition function that measures uncertainty in the Poisson reweighed Laplace learning algorithm (PWLL). In particular, the authors control the exploration versus exploitation tradeoff in the active learning process by introducing a diagonal perturbation in PWLL which produces exponential localization of solutions. The uncertainty sampling-based method [1] queries points near decision boundaries that are most uncertain. Common uncertainty measures include least-confident, smallest margin, and entropy. For comparison, we use the smallest margin acquisition function with Laplace learning as an underlying classifier, which prioritizes points near the decision boundaries. Lastly, we use the seeding method for k-means++ [56] which chooses an initial set of diverse centroids for the k-means clustering algorithm [57] to select centroids within each distinct cluster. One can think of labeling these initial centroid sets to obtain representative and diverse labeled sets within an active learning framework, and we therefore include it as a method for comparison.
For the implementation of MALADY, V-OPT, Random, k-means++, and Unc-Laplace, we utilized the GraphLearning Python package [19] which contains both graph learning and active learning frameworks. For the implementation of Unc. (Norm) [4], we utilize the code obtained from the author’s repository.
@@ 4.3 Hyperparameters Selection
In this section, we outline the selected hyperparameters for our experiments. In particular, the number of nearest neighbors, $N_{n}$, for the weight matrix is a tunable parameter that can be adjusted to control the sparsity of the graph. For the computational experiments, we set the number of steps $s$, auction error tolerance $\epsilon_{min}$, epsilon scaling factor $\alpha$, and the initial epsilon $\epsilon_{0}$ to 100, $1e^{-6}$, 4, and $1e^{-7}$, respectively. We also incorporate class size constraints for the classification which are controlled by the upper bound $U$ and lower bound $B$. In our computational experiments, we use exact class size constraints. Overall, we have outlined all parameters used for the computational experiments in the Supplementary Information file.
@@ 4.4 Performance and Discussion
Information about the data sets used for our computational experiments is detailed in Section 4.1 and Table 1. For each data set, we use five initial labels per class; therefore, the initial labeled set has a total of $5N$ labeled points, where $N$ is the number of classes. Then, we sequentially query 500 additional data elements to label. Moreover, for both hyperspectral datasets, Urban and Jasper Ridge, we use the cosine similarity function (2) for graph weight computation, and we use the Gaussian weight function (1) for the other data sets.
In Figure 4, we show the accuracy performance of each acquisition function averaged over 10 trials, except for the Urban dataset, where we averaged over 5 trials; for all data sets, accuracy was the main evaluation metric. From the results, one can see that our MALADY algorithm has significantly outperformed the other comparison methods in all data sets except COIL-20. For the COIL-20 data set, our MALADY algorithm performs similarly to the best-performing method Unc. Norm [4] which uses a different underlying graph-based classifier than our auction dynamics model.
5 Conclusion
In this paper, we introduce the Multiclass Active Learning with Auction DYnamics on Graphs (MALADY) algorithm which integrates active learning with auction dynamics techniques for semi-supervised learning and data classification and ensures exploitation in the active learning process. The proposed method allows one to obtain accurate results even in the case of very small labeled sets, a common scenario for many applications since it can be costly in time and money to obtain labeled data in many applications. The proposed method leverages a novel acquisition function derived from the dual variables in an auction algorithm. Lastly, the proposed algorithm also allows one to incorporate class size constraints for the data classification task, which improves accuracy even further. Overall, the proposed procedure is a powerful approach for an important machine learning task.